17 HW 13. 10.00 points O you did not Problem 17.9 Dissident stockholder group and cumulative voting|L01721 Ms. is a dissident stockholder. He controls proxies for 32,001 shares. Ms. Ramsey and her friends on the board conteel 52.001 shares O stockholders, whose loyalties are unknown, will be voting the remaining 27.908 shares. The company uses cumulative voting of O Type here to search
px b. How many directors can Ms. Ramsey and her friends be sure of electing? (Do not round intermediate calculations. Round down your answer to the neares c-1. How many directors could Mr. Clark elect if he obtains all the proxies for the votes? (Do not round of d c-2. Will he control the board? O Yes O No e to be elected, and Ms. Ramsey and her friends had 57,001 shares and Mr. Clark had 37,001 shares plus haif the d. If nine directors wer Mr. Clark elect? (Do not ro und intermediate calculations. Round down your answer to the nearest whole number.) of O Type here to search
Answer #1

Given:

Below information is about Midland petroleum:

Number of directors =15

Shares held by Ms Ramsey and her friends on board = 52001 shares

Shares held by others = 27998 shares

Dissident shareholder ( Mr Clark) holds number of proxies = 32001 Shares

Total number of shares outstanding = 52001 + 27998+32001

= 112000 shares

Firm follows cumulative voting.

Answer :

1) No of directors Mr Clark can be sure of electing = 4 directors

No of votes available with Clark =

number of votes equal to his shares * number of seats to be elected.

= 32001*15

= 480015 votes

Percentage of votes = votes with Clark / total votes   

= 480015 / 1680000

= 28.57 %

No of directors that can be elected = 15 * 28.57%

= 4 directors

2) No of votes available with Ramsey and friends = 52001shares * 15

= 780015 votes

Percentage of votes =780015/ 16800000 =

46.43%

Hence no of directors that can be sure = 15* 46.43% = 7 directors

3) on obtaining support of uncommitted shareholders :

Total shares with Clark = 32001 +27998 = 59999 shares

Total percentage of shares = 59999 / 112000 = 53.57%

Hence no of directors that Clark can be sure of having on board = 15*53.57% = 8 directors

Yes he will control the board as out of 15, 8 directors will be his.

4)No of directors to be elected = 9

Shares with Ramsey and her friends = 52001 shares

Shares with Clark = 37001 + (27998/2) = 51000 shares

No of votes with Clark = 51000*9 = 459000 votes

Percentage to total votes = 459000/ (117000*9)

= 43.58%

No of directors he can elect = 9*43.58% = 4 directors

If 9 directors were to be elected and only half of uncommitted shareholders would support, Clark would not get control of the board.